OpenStudy (anonymous):

it takes renee 8 hours to make 7 carvings. at this rate, how many hours will it take him to make 63 carvings? explain how you found the values

OpenStudy (nirmalnema):

renee 8 hours to make 7 carvings for 1 she will take.8/7 hr now for 63 = 63*8/79*8=72 hrs

OpenStudy (anonymous):

thanks

OpenStudy (nirmalnema):

welcome.>!!